OpenStudy (anonymous):

Solve the equation for y. Then find the value of y for the given value of x. 2xy+1=xy+3; x=2

OpenStudy (anonymous):

2xy+1=xy +3 subtract xy from both sides xy+1=3 subtract 1 from both sides xy=2 divide both sides by x y=2/x when x=2 y=2/(2) so y=1

OpenStudy (anonymous):

Thank you!!